OpenStudy (anonymous):

Solve the initial value problem 12(t+1)(dy/dt)−8y=32t, for t>−1 with y(0)=3.

OpenStudy (anonymous):

is this a separable equation?

OpenStudy (loser66):

ODE?

OpenStudy (anonymous):

yes its a differentia eq

OpenStudy (anonymous):

i have an integrating factor of (t+1)^(-2/3) is that correct

OpenStudy (loser66):

integrating factor

OpenStudy (anonymous):

i think there is a simpler way but im not sure

OpenStudy (loser66):

integrating factor is \[\huge e^{\int -\frac{2}{3(t+1)}}\] not just integrand

OpenStudy (anonymous):

that simplifies to (t+1)^(-2/3) though correct?

OpenStudy (loser66):

oh yea, you are right, hehehe. I am sorry.

OpenStudy (anonymous):

okay, so im at a part where i have to integrate \[\frac{ t }{ (t+1)^{5/3} }\]

OpenStudy (anonymous):

how would i do this

OpenStudy (dan815):